Reports 1-1 of 1 Clear search Modify search
MIF (General)
shinji.miyoki - 23:24 Monday 09 March 2020 (13432) Print this report
BS Lateral Position slight adjustment


Waht we did

BS IP lateral position (T) was adjusted from 11um (original position)  to -11 um at IP stage. So, in principle, the only beam spot position should move in the BS surface plane.

Renewed MICH FF was applied but MICH gain is set 0.15.


Related Result or Accidental Result ?

Displacement from 100 Hz to 300 Hz reduced. GR-TR_Y light power increased from 1.14 to 1.20, or error range ??

Binary Range became 594 kpc at most, average 550 kpc.

If we apply more DC gain for MICH loop, a slight reduction might be expected. however, there was other noise sources.


Lock acquire seemed to become difficult. Especially, DC readout pass failed a lot. We need gain adjustment ? or 24.5 Hz peak mitigation ? 

Images attached to this report
Search Help